OpenStudy (anonymous):

Solve, using the product rule, y = (4x-1)(x^2 -5x +1)

OpenStudy (anonymous):

The texts answer given is 12x^2 -38x -1

OpenStudy (aravindg):

ok do u knw hw to use product rule?

OpenStudy (anonymous):

Yes and i've completed the question, but my textbook gives a different answer. I can see where i went wrong, so i just wanted a second opinion :)

OpenStudy (anonymous):

I believe that my textbook has given a wrong answer.

OpenStudy (aravindg):

wow :) dy/dx=4(x^2 -5x +1)+(4x-1)(2x-5) =4x^2-20x+4+8x^2-20x+2x+5 =12x^2-38x+9

OpenStudy (anonymous):

ah i see, my apologies i am an idiot.

OpenStudy (aravindg):

did i make a mistake in constant part or is it ur text ?

OpenStudy (anonymous):

I read t as +38x in my text

OpenStudy (anonymous):

So i didnt see how it would work at all. Thanks very much, I feel like a complete idiot now. Thanks for the help non-the-less. Have a good one mate!!

OpenStudy (aravindg):

yw :)